OpenStudy (anonymous):

I need help please! Ive got most of the question, just not part B! Please help! Will give medal/fan! The table below shows the cube roots of different numbers: Number (x) 8 27 64 125 Cube root (y) 2 3 4 5 Part A: Does the table represent y as a function of x? Justify your answer. Part B: The total cost f(x), in dollars, for renting a bike for x hours is shown below: f(x) = 10 + 20x What is the value of f(100), and what does f(100) represent?

OpenStudy (anonymous):

Is part B just 100 hours of rental?

OpenStudy (catlover5925):

yes and with 100 just plug it in for X

OpenStudy (catlover5925):

10 + 20(100) 20 * 100 = 2000 + 10 = 2010

OpenStudy (catlover5925):

does this make sense???

OpenStudy (anonymous):

Yes, fantastic! Thank you so much!

OpenStudy (catlover5925):

no problem!!

OpenStudy (anonymous):

wow u got it all figured out nice

OpenStudy (anonymous):

great job

OpenStudy (catlover5925):

yep, lots of people tag me in math posts sometimes i get overwhelmed, lol

OpenStudy (anonymous):

thx for the medal Shiva_

OpenStudy (anonymous):

u r like a human calculator lol

OpenStudy (anonymous):

im sorry but this is way way off topic i dont mean to be weird but gosh you r just so Beautiful @Shiva_

OpenStudy (anonymous):

Haha! I wish I could take that complement! The girl in the picture is actually Marine Lorphelin. :P Sorry! haha!

OpenStudy (anonymous):

Well dang she is still pretty tho

OpenStudy (anonymous):

Haha Yes! She is actually Miss France 2013 :)

OpenStudy (catlover5925):

@jackson21 btw i gave you the medal

OpenStudy (anonymous):

sorry well thanks @Catlover5925 P.S. ur really smart

OpenStudy (catlover5925):

thanks!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

that profile is kinda cute